A Gallup poll of 1,236 adults showed that 12% of the respondents believe that it is bad luck to walk under a ladder. SupposeSuppose a certain experiment yields a binomial distribution with n = 64 trials and a probability of success of p = 0.409. WhaA pharmaceutical company receives large shipments of aspirin tablets. The acceptance sampling plan for each shipment is to raAccording to a recent poll, 39% of U.S. voters approve of the Presidents job performance. Suppose that 50 U.S. voters are seBased on the results of the previous problem, what numbers of voters who approve of the Presidents job performance would be
Answer #1

1) Answer is

Although selections are not independent , they can be treated as if they are independent using bu applying the 5% rule.

Note : Since selections are done without replacement , each selection is dependent on the earlier selections , so they are not independent . But sample size is musch less than 5% of population size (5% of1236 = 61.8 ) , thus they can be treated as indpendent.

2) Let X follow Binomial with n= 64 , p =0.409

Then probability mass function of X is

P(X=x)= \binom{64}{x}0.409^{x}(1-0.409)^{64-x}, x=0,1,2,....64

To find

P(X\leq 27)=\sum_{x=0}^{27} \binom{64}{x}0.409^{x}(1-0.409)^{64-x}

= 0.634

3) Let X be the number of tablets which do not meet specification

X follow Binomial with n= 58 , p =0.04

To find

P(X\leq 1)=\sum_{x=0}^{1} \binom{58}{x}0.04^{x}(1-0.04)^{58-x}

= \binom{58}{0}0.04^{0}(1-0.04)^{58-0}+\binom{58}{1}0.04^{1}(1-0.04)^{58-1}

= 0.320

4) Let X be the number of voters who approve of Presidents job performance

X follow Binomial with n= 50 , p =0.39

We know that for Binomial distribution

mean = np = 50*0.39 = 19.5

variance = np(1-p) = 50*0.39*(1-0.39) = 11.9

Standard deviation = \sqrt{np(1-p)}=3.4

5) As n is large ( n> 50) and np = 19.5 > 10 and n(1-p) = 30.5 > 10 , we can use Normal approximation to Binomial .

Using empirical probability method , we know that 95% of the observation lie between 2 standard deviation from mean .

Outside the interval, (mean -2 SD , mean +2 SD) , values are considered unusual

Thus minimum value = mean - 2 SD

= 19.5- 2 * 3.4 = 12.7 = 13 ( nearest interger)

maximum value = mean + 2 SD

= 19.5 + 2 * 3.4 = 26.3 =26 ( nearest interger )

Out of 50 voters surveyed , 13 or fewer voters would be significantly low and 26 or more would be significantly high .